OpenStudy (anonymous):

What is the square root of r^5 ???

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large \sqrt{r^5} = (r^5)^{1 \over 2} = r^{2.5}}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow r^{2.5} = r^2 \times r^{0.5} = r^2 \times \sqrt r} = r^2\sqrt r\)

Parth (parthkohli):

@cornitodisc the question is different :)

Parth (parthkohli):

Another way to think about this is: \( \color{Black}{\Rightarrow \sqrt{r^5} = \sqrt{r \times r \times r \times r \times r} = r^2\sqrt r }\)

Parth (parthkohli):

@LilGorgeousTy get it?

OpenStudy (anonymous):

\[\huge \color{green}{\sqrt{r^5} = \sqrt{r \times r \times r \times r \times r} = r \times r \sqrt{r}}\] \[\huge \color{green}{ = r^2\sqrt{r}}\]

OpenStudy (anonymous):

Yeahhh. I was just making sure if thats what i had too.. Thankss (:

Parth (parthkohli):

You're welcome!

OpenStudy (anonymous):

\[\sqrt{r^5}=\sqrt{r^2}.\sqrt{r^2}.\sqrt{r}=r.r.\sqrt{r}=r^2\sqrt{r}\]

Parth (parthkohli):

@cornitodisc Yes!

Parth (parthkohli):

Use \cdot for multiplication :)

OpenStudy (anonymous):

@ ParthKohli... :)